please help me solve this problem .
A ball is dropped from a height of 6 meters and returns to about 7/8 of its
previous height on each bounce. About how far will the ball travel before it
comes to rest?

Answer by Boreal(15235)

This is a geometric series with 4=7/8
Sum is a/(1-r)=6/(1-(7/8))=6/(1/8)=48 m.